OpenStudy (anonymous):

ok lets see who can help me!!! what times what gives u -20 but when u add those two numbers up u get -5??? example: -5 times -3=15 and -5+-3=8

OpenStudy (anonymous):

x = 1/2 (5-sqrt(105)) x = 1/2 (5+sqrt(105))

OpenStudy (anonymous):

ok